OpenStudy (anonymous):

What is the solution to the equation 9^(x + 2) = 27 ?

OpenStudy (anonymous):

x=-1/2

OpenStudy (anonymous):

THANKS !

OpenStudy (anonymous):

Sure =)

OpenStudy (amistre64):

3^(2(x+2)) = 3^3 2(x+2) = 3 x+2 = 3/2 x = 3/2 - 2 = -1/2 yep, that does work out